A 47.7 g sample of Cu is added to 116 mL of 5.5 M HNO3. Assuming the following reaction is only one that occuts;




Cu(s) + HNO3 ----> Cu(NO3)2 (aq) + H2O(ı) + NO(g)




Will the Cu react completely? What is the limiting reagent and what is the remaining compound in grams? How many grams of copper(II) nitrate will be obtained?

1
Expert's answer
2021-11-08T08:11:20-0500

3 Cu (s) + 8 HNO3 (aq) → 3 Cu(NO3)2 (aq) + 2 NO (g) + 4 H2O (l)


Moles of Cu"=\frac{47.7}
{63.546}=0.75moles"


Moles of HNO3 "=5.5×\frac{116}{1000}=0.638"



No


"HNO_3" Is the limiting reagent



Moles of Cu(NO3)2"=0.638×\frac{3}{8}=0.239moles"


Cu(NO3)2 Molar mass: 187.56 g/mol


Mass "0.239\u00d7187.56=44.83grams"